Can someone please show the steps for this and give the ordered pair? use the method of substitution to solve the system of linear equations.
2x-y=4
7x-3y=16

Answer by fcabanski(1391)

Multiply the first equation by 3: 3*(2x-y=4) = (6x-3y=12)


Subtract that equation from the second equation: 7x-3y=16 - (6x-3y=12) becomes 7x-6x -3y + 3y = 16-12 ---> x=4


Substitute that into either equation (try the first): 2(4) - y = 4 ---> 8 - y = 4 ---> 4=y


The ordered pair solution is (4,4)

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)